Observing Log for PISCES night of 29 February 2012 UT

AZ / OSURC mini-queue with Hill et al., Close et al., McGreer et al., Woodward et al.,

Observer: JHill (LBTO) wearing his AZ observer hat
AO & Observation Support: DMiller (LBTO), JCGuerra (LBTO)
Telescope Operator: SAllanson (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: TShih (LBTO)

Summary

Sky is photometric, but seeing is 2 arcsec on the guider (and 4 arcsec on the DIMM) so not the best AO conditions. Also serious problems with the W-unit camera lens temperature being too cold for the first half. We worked around that by raising the instrument cooling temperature and having warmer weather. We spend the first half of the night ironing wrinkles out of other things and taking photometric calibrations in bad seeing.

Once the camera lens was warm enough, then we ran into other problems getting AcquireRef to work properly. One problem was that the rerotator was not operating properly. A second problem was that the loop would not close in a stable fashion. Alfio suggests that this might be another symptom of the camera lens BCU being still too cold at 6.5 degC.

We finally got the loop to close nicely at 11:37 UT. We spent the last hour of the night learning the proper way to observe the comet nucleus in pretty good conditions with low-Strehl in J-band (numbers to follow later). We did get a little bit of imaging data on the comet in closed loop.

Perhaps the best part of the night was that Walter was able to get parallel sky time to align the LUCI1 pupil.

Data Summary for Tonight

  • open loop standard stars in several filters
  • a small amount of data on C/2009 P1 Garradd for Woodward (~0:40)
  • sky flats in the morning
  • darks in the morning

Bad Images

Images 10020, 10047, 10230, 10385 appear to be all zeros on ds9. However the corresponding raw images w/o full headers seem to be OK. See Issue 3980

Preparations

Jeff has cleared the rime ice off of the weather station and shutter door seals (1.5 hours of work on the roof).

00:20 UT Sky is mostly clear, but humidity is too high (96%) to open. Thus, we were not able to open early for sky flats.

00:21 UT Test GUIDE ACE preset from IDL succeeds up to where it could not find a star.

00:23 UT Test TRACK ACE preset also succeeds (such as a closed dome preset does).

00:25 UT Park at zenith for some PISCES web camera work (replugging USB connectors). (10 minutes)

01:05 Moved all the 2011 ephemerides into a subdirectory (/lbt/ephemeris/2011ephemerides/) to clean up the choose box for the NSIGUI.

01:14 Created 4 nights worth of ephemerides for 2009P1.

Details

01:15 Humidity is down to 78% so it looks like we are going to get to open. Side Vent Doors are open as we ventilate the mirrors.

For most of the night, JCG is driving the AO system and PISCES, DLM is issueing the observing scripts, and JMH is recording in the wiki.

Open

01:36 UT Open T=-8.1 degC D=-15 degC and dropping. Steady Wind 8 m/sec from SW. Photometric although some cirrus visible below EL=2 deg in SW. Primary Mirrors are ~5 degC too warm (from being closed a couple nights and a cold front going through last night).

01:45 ACQUIRE Preset to a pointing star WT10_252, failed for shell not set.

01:47 Preset successful

(suggestion: make Paddle GUI for pointing alignment at non-zero parallactic angle - Steve is thinking about it)

01:50 Doug restarts PISCES server - filter wheel numbering is off.

01:55 ACTIVE preset to collimate - but guiding failed. ?? bad acq image ??

01:57 ACTIVE preset succeeds - initial seeing on the guider is 2.0 arcsec FWHM - primary mirrors are now 4 degC above ambient

(suggestion: make automated GS selection for PISCES rotator angle -- everybody knows this but we don't have it implemented yet)

AO observations are running from directory /OldRepository/PISCES_Data/Feb12_29/


02:02 Begin overhead for Close et al. (I'm trying in this log to keep track of both hours of overhead and hours spent integrating on each of the projects.)

02:10 Preset to HIP43121 in regular ACE mode PA=290 (not ADI, to make sure that all systems are working)

Why is the star spiraling in the guide box??? Perhaps PCS is confused?

02:13 Stop and restart PCS (5 minutes lost)

02:15 Resend the preset - forget to reset IE/CA

02:16 Resend the preset - succeeds at the level of guiding although seeing image motion is jumping all around -- failed because of camera lens temperature.

02:23 Resent the preset star disappears on ACQ image after one frame

02:24 Resent the preset star disappears on ACQ image after one frame

02:27 GUIDE preset fails for guiding not converging

02:27 ADAPTIVE ACE

Changed Guide Correction Factor from 0.8 to 0.5 for RFBG - We were going to make exactly this change during next week's technical block in any case.

Switch to a target a higher elevation so that we suffer less with the seeing.


02:36 Stop Clock for HIP43121 overhead -- Start Clock for DGTau overhead

DIMM reports 4 arcsec seeing (not sure if that means anything or not)

02:37 Stop and Restart GCSR (probably it picked up the Guide Correction Factor automatically)

02:38 GUIDE preset on-axis, since DGTau doesn't have useful guide stars. There was not any junk image displayed this time, on the GCSGUI so the acquisition and stabilization of the guide star was completely normal.

02:44 ACTIVE preset on-axis to DGTau (waiting to fix the astigmatism) -- image didn't arrive within timeout

02:46 Stop GCSR, reboot the azcamserver machine, restart GCSR

02:49 resend the ACTIVE preset - succeeds - seeing on guider is 1.8-2.2 arcsec at EL=68. Yuk.

02:54 Try ADAPTIVE ACE TRACK preset to DGTau - but Doug finds a script problem in IDL. We stop to repair it since the seeing is 2 arcsec and the camera lens is cold.

03:04 Send ADAPTIVE ACE TRACK preset to DGTau - succeeds but we don't see a star on the CCD39. Why? Seems to be that pointing needs updated -- telescope is cooling fairly rapidly.

03:09 ACQUIRE preset to DGTau - reset IE/CA

03:14 ADAPTIVE ACE TRACK fails because Camera Lens Temperature too low 5.7 degC. See Issue 3967 (Ambient is -6.2 degC)

03:26 Turn W unit off and on to see if this helps the camera lens problems.

03:35 ADAPTIVE ACE TRACK preset -- succeeded, and we see the star.

03:38 RunAO fails because Camera Lens too cold (T=4.5)

03:40 Wind is increasing. T=-5.4 degC D=-21 degC Sky is still clear.


03:43 start overhead for C/2009 P1 Garradd -- Doug and John are slowly trying out all the observing modes and coordinates for the week while Juan Carlos is chewing on the camera lens temperaure problem. Oops, the target is too low in the north for AO - this is a second half target. NSIGUI crashed.


Checking for Condensation on PISCES

03:44 start overhead for HIP43121 - Back to HIP43121 to check PISCES window in ACTIVE mode - humidity was high over the previous 24 hours, although dome was warm.

03:52 ACTIVE preset to HIP43121 at PA=-77

03:58 Steve sending DETXY offsets from IRAF LBTtools to move the star around the PISCES image. (AO loop is OPEN and there is bad seeing.) These images start with pisces.20120228.0029.fits and cover 29-34.

04:04 Doug resets the archive parameters so we now have images with tonight's UT date - restarting the numbers at 1.

Calibrating the ND Filter

04:07 Move to zenith to put in the ND filter. We are going to measure the "photometry" and the "vignetting" of the ND filter in front of PISCES.

04:20 Found an alarm going off in the LUCI1 rack on the telescope.

04:35 LUCI alarm has been reset by Tim.

04:43 ACTIVE preset but pointing is fairly far off.

04:44 ACQUIRE preset to HIP43121 to adjust pointing (off by ~15 arcsec)

04:46 ACTIVE preset to HIP43121

04:48 Primary mirrors are now within ~1 degC of chamber ambient, and telescope steel is about 2 degC too warm. Seeing on guider is now sub-2 arcsec (but only barely) at EL=67. Outside ambient is -5.8 degC

04:50 Steve sending offsets to move the image around on PISCES. These data are PISCES images 002-009. JC says we have some vignetting in the lower part of the chip.

05:02 Move to zenith to remove the ND filter.


Observing Standard Stars

Sky is still photometric although the wind could be kicking up some dust.

05:19 ACTIVE preset to BS9142 at PA=190 (-170) starting at EL=70 - succeeded. bs9142.obs

Seeing is bouncing between 2 and 3 arcsec on the guider - even worse than earlier.

(suggestion: allow the "header" command to be somewhere else than position 1 in the script so that the script can be reentrant after the preset and still get header info. OR read the header in position 1 all the time.)

05:42 Unrestrict DD variables because of "LOOP CLOSED"

05:43 observe, BS9142_test.obs

Doing dice 5 dithers on filters 4 J, 5 Ks, 6 H 15 sec each starting with image 10017 (exposures before that were JC's exposure time checks I think) Image 10026 is bad. This group of images through 10031.

06:13 observe, /OPEN, BS9142_test.obs (now with different filters and exptimes)

Filters 1 H2, 7 Br-g, 8 FeII with exposure=60sec starting with image 10036. Image 10047 is bad. Last image is 10050.


06:42 ADAPTIVE ACE ACTIVE preset to BS9142 PA=-220

06:43 RunAO fails with temperature = 5.1

06:52 Guider seeing has improved modestly to 1.8 arcsec, while DIMM is reporting 2.2.

06:57 Call Armando in Italy hoping that he knows Alfio's home/cell phone number, but Armando didn't answer.

07:02 Increase the Right Instrument Cooling Offset from -1.0 to +1.0 in hopes of warming the W-unit by a couple degrees and by heat transfer the BCU that controls the camera lens. This increased the Right IC Setpoint from -4.97 to -2.97 degC. "BCU47" is 5.4 degrees.

Object: BS9142
Dither
X Y frames Exposure Filter PISECS Image
0.00000 0.00000    
" " 1 of 1 15.0000 J pisces.20120229.10017.fits STANDARD
5.00000 5.00000    
" " 1 of 1 15.0000 J pisces.20120229.10018.fits STANDARD
-5.00000 5.00000    
" " 1 of 1 15.0000 J pisces.20120229.10019.fits STANDARD
-5.00000 -5.00000    
" " 1 of 1 15.0000 J pisces.20120229.10020.fits STANDARD
5.00000 -5.00000    
" " 1 of 1 15.0000 J pisces.20120229.10021.fits STANDARD
Dither
X Y frames Exposure Filter PISECS Image
0.00000 0.00000    
" " 1 of 1 15.0000 Ks pisces.20120229.10022.fits STANDARD
5.00000 5.00000    
" " 1 of 1 15.0000 Ks pisces.20120229.10023.fits STANDARD
-5.00000 5.00000    
" " 1 of 1 15.0000 Ks pisces.20120229.10024.fits STANDARD
-5.00000 -5.00000    
" " 1 of 1 15.0000 Ks pisces.20120229.10025.fits STANDARD
5.00000 -5.00000    
" " 1 of 1 15.0000 Ks pisces.20120229.10026.fits STANDARD
Dither
X Y frames Exposure Filter PISECS Image
0.00000 0.00000    
" " 1 of 1 10.0000 H pisces.20120229.10027.fits STANDARD
5.00000 5.00000    
" " 1 of 1 10.0000 H pisces.20120229.10028.fits STANDARD
-5.00000 5.00000    
" " 1 of 1 10.0000 H pisces.20120229.10029.fits STANDARD
-5.00000 -5.00000    
" " 1 of 1 10.0000 H pisces.20120229.10030.fits STANDARD
5.00000 -5.00000    
" " 1 of 1 10.0000 H pisces.20120229.10031.fits STANDARD
Dither
X Y frames Exposure Filter PISECS Image
0.00000 0.00000    
" " 1 of 1 60.0000 Br-g pisces.20120229.10036.fits STANDARD
5.00000 5.00000    
" " 1 of 1 60.0000 Br-g pisces.20120229.10037.fits STANDARD
-5.00000 5.00000    
" " 1 of 1 60.0000 Br-g pisces.20120229.10038.fits STANDARD
" " 1 of 1 60.0000 Br-g pisces.20120229.10039.fits STANDARD
5.00000 -5.00000    
" " 1 of 1 60.0000 Br-g pisces.20120229.10040.fits STANDARD
Dither
X Y frames Exposure Filter PISECS Image
0.00000 0.00000    
" " 1 of 1 60.0000 FeII pisces.20120229.10041.fits STANDARD
5.00000 5.00000    
" " 1 of 1 60.0000 FeII pisces.20120229.10042.fits STANDARD
-5.00000 5.00000    
" " 1 of 1 60.0000 FeII pisces.20120229.10043.fits STANDARD
-5.00000 -5.00000    
" " 1 of 1 60.0000 FeII pisces.20120229.10044.fits STANDARD
5.00000 -5.00000    
" " 1 of 1 60.0000 FeII pisces.20120229.10045.fits STANDARD
Dither
X Y frames Exposure Filter PISECS Image
0.00000 0.00000    
" " 1 of 1 60.0000 H2 pisces.20120229.10046.fits STANDARD
5.00000 5.00000    
" " 1 of 1 60.0000 H2 pisces.20120229.10047.fits STANDARD
-5.00000 5.00000    
" " 1 of 1 60.0000 H2 pisces.20120229.10048.fits STANDARD
-5.00000 -5.00000    
" " 1 of 1 60.0000 H2 pisces.20120229.10049.fits STANDARD
5.00000 -5.00000    
" " 1 of 1 60.0000 H2 pisces.20120229.10050.fits STANDARD


Trying the Comet again

07:15 ACTIVE preset M5_0619 gs=5 PA= to EL=31.1 - succeeds but is rather low

07:26 ACTIVE preset WT10_387 gs=1 PA= to EL=37.5 - but we seem to have found the wrong guide star due to pointing drift.

07:29 ACQUIRE preset WT10_387 gs=0

In the meantime we think the camera lens BCU might be warm enough.

07:32 NSI ADAPTIVE ACE TRACK fails because we can't hijack the reference star - need to think more about that later.

Q1135-0011

07:40 ADAPTIVE ACE ACTIVE preset for Q1135-0011 gs=4 rs=1 PA= failed because NSI flag was still on.

07:45 ADAPTIVE ACE ACTIVE preset for Q1135-0011 gs=4 rs=1 PA=-50 - succeeded with Z6

Seeing on the guider has improved to 1.3 arcsec!

07:52 Reduce to RIGHT IC offset to 0.0 as the camera lens BCU is up to 7.3.

07:52 AcquireRef failed, but we don't understand why -- it seems that W is measuring the star 2.5 mags too faint.

07:53 resend the preset, but the loop was not able to successfully close.

08:00 ACQUIRE for Q1135-0011 on the reference star gs=1 ts=1 PA=-50

08:01 Reduce the Right IC offset to -1 degC (its original value)

08:06 ADAPTIVE ACE ACTIVE

08:14 DX Primary Mirror Panic (probably because AO offload some garbage forces)

AO loses contact with anemometer and sets TSS. Restart AO arbitrator. Clear Active Optics on M1. Raise M1.

08:20 Continue having trouble with this 14.5 magnitude star, so decide to try a brighter (non-science) object.

AO0423

08:29 ADAPTIVE ACE ACTIVE preset to AO423 gs=1 PA=-70 - Seeing varies from 1-1.5 arcsec at EL=75.8

08:33 RunAO failed - it bent the secondary into a crazy shape (wild quadrafoil?) - a similar bad thing like it was doing on the fainter star.

08:35 JC tries an AcquireRef through the engineering interface - it fails in the same way.

08:43 Reset and reflatten the shell. Clear C00 on DX primary.

08:45 ACTIVE preset to AO423 - AcquireRef fails the same way

08:53 Reflatten the shell. Clear C00 on DX primary

08:54 in parallel JC is power cycling the W-unit and restarting the W software.

08:55 ACTIVE preset to AO423

09:01 ADAPTIVE ACE ACTIVE preset to AO423 - seeing on the guider goes below 1 arcsec for the first time tonight. Preset failed for "Lamp is not ready".

09:03 Resend - fails same way. Turn on flower pot.

09:04 Resend - preset succeeds.

09:06 RunAO - fails with lots of quadrafoil. Suspect that the rerotator is not going to the proper angle (or something like that).

09:09 Rest shell, stop and restart AOSR. Clear C00 (maybe not in that order). Flatten the shell.

09:14 ADAPTIVE ACE ACTIVE preset to AO423

09:16 RunAO - but rerotator is still at 0.000 degrees.

09:20 preset failed - reading shared buffer masterdiagnostic

09:22 Press "Operate"

09:25 Wind is picking up, so we need to find another target in the East.

AO499

09:26 restart the W unit again, making more effort to leave it in the default configuration.

09:30 ACTIVE preset AO499 gs=2 PA=-90 - succeeded at EL=65 seeing on the guider is 0.8 - 0.9 arcsec FWHM.

09:32 ADAPTIVE ACE ACTIVE preset (but he sent it to the wrong target)

09:34 ADAPTIVE ACE ACTIVE preset AO499 - it got confused about 2 PresetAOs running

09:37 ADAPTIVE ACE ACTIVE preset AO499 - succeeded

09:40 Restart the Arbitrator, and then powered on the W-unit from the AOSGUI

09:44 ADAPTIVE ACE ACTIVE preset AO499

09:46 RunAO - failed, but for a new reason - Timeout error from autogain. The good news here is that the rerotator is now working properly.

WFS still thinks the loop is closed.

09:50 RunAO with flag - illegal command for state operational

09:51 StopAO - illegal command for state operational

09:52 RunAO - illegal command for state operational

09:53 ADAPTIVE ACE ACTIVE preset AO499 -

09:56 RunAO - completed successfully, but then the loop immediately opened because of skip frames. Why?

10:00 ADAPTIVE ACE ACTIVE preset AO499 - rerotator complaining of not ready, but that may be normal ??

10:02 RunAO - succeeded but loop opened.

10:04 Turn off flower pot.

10:05 ADAPTIVE ACE ACTIVE preset AO499 - succeeded - guider 0.9 DIMM 1.2

10:06 RunAO - succeeded but loop opened.

10:07 Turn off both primary mirror ventilations

10:12 JC tries to close the loop with manually applied gains (but not exactly at this time)

10:14 ADAPTIVE ACE ACTIVE preset AO499

10:16 RunAO - succeed but loop opened

Object: AO499
Dither
X Y frames Exposure Filter PISECS Image
0.00000 0.00000    
" " 1 of 1 0.800000 Ks pisces.20120229.10052.fits SCIENCE
5.00000 5.00000    
" " 1 of 1 0.800000 Ks pisces.20120229.10053.fits SCIENCE
-5.00000 5.00000    
" " 1 of 1 0.800000 Ks pisces.20120229.10054.fits SCIENCE
-5.00000 -5.00000    
" " 1 of 1 0.800000 Ks pisces.20120229.10055.fits SCIENCE
5.00000 -5.00000    
" " 1 of 1 0.800000 Ks pisces.20120229.10056.fits SCIENCE

Power cycle AO system

10:25 Decide to power cycle and restart entire AO system. We also need to power off and restart Doug.

10:29 reinitialize hexapod

Of course, now the camera lens is down to 5.7 degC.

10:38 ADAPTIVE ACE ACTIVE preset AO499 failed in presetAO with "Arbitrator interface generic error"

10:42 ADAPTIVE ACE ACTIVE preset AO499 PA=-40 - succeeded.

10:49 camera lens at 6.0 deg, so now RunAO - rerotator is NOT updating AGAIN!!!!!!!!!! AaargHHH! JC restarted the arbitrator before the AcquireRefAO timed out so this didn't end cleanly even though the failure was obvious.

10:54 Clear C00

11:00 Rest shell. Stop and restart AOSR. Set shell.

11:10 Doug restarting everything again. We remain unable to contact Alfio or Simone.

11:18 ADAPTIVE ACE ACTIVE preset AO499 PA=-40 but pointing getting close to the edge.

11:21 RunAO - fails - rerotator is still not updating.

11:24 Trying another restart sequence of something suggested by Enrico. And it started working, but none of us understand how this related to Enrico's restarting of WFS arbitrator and Rerotator process.

11:29 ADAPTIVE ACE ACTIVE preset AO499 PA=-40 - succeeded

11:30 RunAO - to quote Steve "the rerotator problem is resolved, but the AquireRef autogain is still broken"

Turn off flower pot now.

11:35 ADAPTIVE ACE ACTIVE preset AO499 PA=-20 - succeeded

11:37 RunAO -- InformationTechnology WORKS! Loop is closed ---- by the sheer presence on the network of Alfio!!!!

Alfio suggests that it could be that the temperature of the camera lens BCU really needs to be 8 degC rather than the 6.8 degC that we have now.

Less seriously, John thinks that some critical part of the AO software only runs on Alfio's laptop.

11:44 Adjust Right IC Offset to +1 to warm the BCU47 some more.

11:48 Doug taking some dithered data on AO499 (whatever bright star it is -- BD+16 2612) At last we have a nice looking diffraction pattern.

11:54 JC taking some optical loop data


Comet 2009 P1 again

11:57 ACQUIRE preset for WT10_387 PA=-210 now at EL=62

BCU47 up to 7.3 degC. Cooling water is -0.68 degC.

12:01 ACTIVE preset for WT10_387 to collimate - seeing is 0.7 arcsec on guider, 1.0 arcsec on DIMM (this guide star was a binary so not good coordinates)

12:20 NSI ADAPTIVE ACE TRACK preset to the comet with a faked up reference star coordinate and it almost worked except that the comet had moved by the time PresetAO moved the stages to the reference star position.

12:24 NSI ADAPTIVE ACE TRACK preset .......

12:26 move stage coordinates to 0,0

12:27 RunAO - no star found in TV frame

12:28 NSI ADAPTIVE ACE TRACK preset ..... (but he forgot to set NSI on this one)

12:30 move stage to center the comet (as the reference star) but couldn't find it

12:33 NSI ADAPTIVE ACE TRACK preset .....

12:33 move stage to center the comet (as the reference star) - the nucleus looks nice and crisp - measured at 14.3 magnitude (we had guessed 11)

12:36 NSI ADAPTIVE ACE TRACK preset .....

12:37 move stage to center the comet (as the reference star) now with less filter in the acq image

12:39 RunAO -- loop closed!!!!! on a comet nucleus

12:44 Test image with PISCES

12:45 Doug starts dither sequence with single 15 sec exposures - offset failed because the offset moved the stage back to the original position (the inverse of our manual motion). So we need to do it without manual intervention.

See Issue 3970 about the need for a Non-sidereal Reference Star in Adaptive Presets.

12:50 NSI Adaptive ACE TRACK preset

12:51 RunAO (recentered on its own) 66 modes TT 6.0

12:55 Doug starts dither sequence, but offset pointing 0,0 in detxy caused the loop to open. Why? Is this the same bug that caused LBC binocular offsets to fail in BP14. Or a new and unique bug?

12:58 preset

12:59 RunAO

13:04 Doug tries another scheme of dithering with the loop closed.

10078 was the center image.

13:15 repeat the dithers in Ks filter
First Ks image is 10084.
Second Ks image is 10085
10086 has the loop opened
(there was also a human error in the dither pattern)

13:21 AO loop opens - presumably because the sky is getting too bright.

" " 1 of 1 15.0000 J pisces.20120229.10078.fits SCIENCE
" " 1 of 1 15.0000 J pisces.20120229.10080.fits SCIENCE
" " 1 of 1 15.0000 J pisces.20120229.10081.fits SCIENCE
" " 1 of 1 15.0000 J pisces.20120229.10082.fits SCIENCE
" " 1 of 1 15.0000 J pisces.20120229.10083.fits SCIENCE
" " 1 of 1 15.0000 Ks pisces.20120229.10084.fits SCIENCE
" " 1 of 1 15.0000 Ks pisces.20120229.10085.fits SCIENCE

Sorry, the sun is rising, so no standard for this set of observations.

13:22 change IE by 200 arcsec to offset for some sky frames

First Ks background is 10087

13:27 start J background frames - with rising twilight

Last J background is 10096

" " 1 of 5 15.0000 Ks pisces.20120229.10087.fits BACKGROUND
" " 2 of 5 15.0000 Ks pisces.20120229.10088.fits BACKGROUND
" " 3 of 5 15.0000 Ks pisces.20120229.10089.fits BACKGROUND
" " 4 of 5 15.0000 Ks pisces.20120229.10090.fits BACKGROUND
" " 5 of 5 15.0000 Ks pisces.20120229.10091.fits BACKGROUND
" " 1 of 5 15.0000 J pisces.20120229.10092.fits BACKGROUND
" " 2 of 5 15.0000 J pisces.20120229.10093.fits BACKGROUND
" " 3 of 5 15.0000 J pisces.20120229.10094.fits BACKGROUND
" " 4 of 5 15.0000 J pisces.20120229.10095.fits BACKGROUND
" " 5 of 5 15.0000 J pisces.20120229.10096.fits BACKGROUND

Sky flats

13:29 Move to zenith facing west (280) to park telescope for sky flats

sky flat directory

*/OldRepository/PISCES_Data/Feb12_Calib/*

images 160-163 were wrong filter

beware of sub-directories of sub-directories

DARKS

DARK directory

*/OldRepository/PISCES_Data/Feb12_Calib/DARKS*

" " 1 of 20 1.00000 blank/dark pisces.20120229.10252.fits DARK
" " 2 of 20 1.00000 blank/dark pisces.20120229.10255.fits DARK
" " 3 of 20 1.00000 blank/dark pisces.20120229.10256.fits DARK
" " 4 of 20 1.00000 blank/dark pisces.20120229.10257.fits DARK
" " 5 of 20 1.00000 blank/dark pisces.20120229.10258.fits DARK
" " 6 of 20 1.00000 blank/dark pisces.20120229.10259.fits DARK
" " 7 of 20 1.00000 blank/dark pisces.20120229.10260.fits DARK
" " 8 of 20 1.00000 blank/dark pisces.20120229.10261.fits DARK
" " 9 of 20 1.00000 blank/dark pisces.20120229.10262.fits DARK
" " 10 of 20 1.00000 blank/dark pisces.20120229.10263.fits DARK
" " 11 of 20 1.00000 blank/dark pisces.20120229.10264.fits DARK
" " 12 of 20 1.00000 blank/dark pisces.20120229.10265.fits DARK
" " 13 of 20 1.00000 blank/dark pisces.20120229.10266.fits DARK
" " 14 of 20 1.00000 blank/dark pisces.20120229.10267.fits DARK
" " 15 of 20 1.00000 blank/dark pisces.20120229.10268.fits DARK
" " 16 of 20 1.00000 blank/dark pisces.20120229.10269.fits DARK
" " 17 of 20 1.00000 blank/dark pisces.20120229.10270.fits DARK
" " 18 of 20 1.00000 blank/dark pisces.20120229.10271.fits DARK
" " 1 of 20 3.00000 blank/dark pisces.20120229.10272.fits DARK
" " 2 of 20 3.00000 blank/dark pisces.20120229.10273.fits DARK
" " 3 of 20 3.00000 blank/dark pisces.20120229.10274.fits DARK
" " 4 of 20 3.00000 blank/dark pisces.20120229.10275.fits DARK
" " 5 of 20 3.00000 blank/dark pisces.20120229.10276.fits DARK
" " 6 of 20 3.00000 blank/dark pisces.20120229.10277.fits DARK
" " 7 of 20 3.00000 blank/dark pisces.20120229.10278.fits DARK
" " 8 of 20 3.00000 blank/dark pisces.20120229.10279.fits DARK
" " 9 of 20 3.00000 blank/dark pisces.20120229.10280.fits DARK
" " 10 of 20 3.00000 blank/dark pisces.20120229.10281.fits DARK
" " 11 of 20 3.00000 blank/dark pisces.20120229.10282.fits DARK
" " 12 of 20 3.00000 blank/dark pisces.20120229.10283.fits DARK
" " 13 of 20 3.00000 blank/dark pisces.20120229.10284.fits DARK
" " 14 of 20 3.00000 blank/dark pisces.20120229.10285.fits DARK
" " 15 of 20 3.00000 blank/dark pisces.20120229.10286.fits DARK
" " 16 of 20 3.00000 blank/dark pisces.20120229.10287.fits DARK
" " 17 of 20 3.00000 blank/dark pisces.20120229.10288.fits DARK
" " 18 of 20 3.00000 blank/dark pisces.20120229.10289.fits DARK
" " 19 of 20 3.00000 blank/dark pisces.20120229.10290.fits DARK
" " 20 of 20 3.00000 blank/dark pisces.20120229.10291.fits DARK
" " 1 of 20 5.00000 blank/dark pisces.20120229.10292.fits DARK
" " 2 of 20 5.00000 blank/dark pisces.20120229.10293.fits DARK
" " 3 of 20 5.00000 blank/dark pisces.20120229.10294.fits DARK
" " 4 of 20 5.00000 blank/dark pisces.20120229.10295.fits DARK
" " 5 of 20 5.00000 blank/dark pisces.20120229.10296.fits DARK
" " 6 of 20 5.00000 blank/dark pisces.20120229.10297.fits DARK
" " 7 of 20 5.00000 blank/dark pisces.20120229.10298.fits DARK
" " 8 of 20 5.00000 blank/dark pisces.20120229.10299.fits DARK
" " 9 of 20 5.00000 blank/dark pisces.20120229.10300.fits DARK
" " 10 of 20 5.00000 blank/dark pisces.20120229.10301.fits DARK
" " 11 of 20 5.00000 blank/dark pisces.20120229.10302.fits DARK
" " 12 of 20 5.00000 blank/dark pisces.20120229.10303.fits DARK
" " 13 of 20 5.00000 blank/dark pisces.20120229.10304.fits DARK
" " 14 of 20 5.00000 blank/dark pisces.20120229.10305.fits DARK
" " 15 of 20 5.00000 blank/dark pisces.20120229.10306.fits DARK
" " 16 of 20 5.00000 blank/dark pisces.20120229.10307.fits DARK
" " 17 of 20 5.00000 blank/dark pisces.20120229.10308.fits DARK
" " 18 of 20 5.00000 blank/dark pisces.20120229.10309.fits DARK
" " 19 of 20 5.00000 blank/dark pisces.20120229.10310.fits DARK
" " 20 of 20 5.00000 blank/dark pisces.20120229.10311.fits DARK
" " 1 of 20 10.0000 blank/dark pisces.20120229.10312.fits DARK
" " 2 of 20 10.0000 blank/dark pisces.20120229.10313.fits DARK
" " 3 of 20 10.0000 blank/dark pisces.20120229.10314.fits DARK
" " 4 of 20 10.0000 blank/dark pisces.20120229.10315.fits DARK
" " 5 of 20 10.0000 blank/dark pisces.20120229.10316.fits DARK
" " 6 of 20 10.0000 blank/dark pisces.20120229.10317.fits DARK
" " 7 of 20 10.0000 blank/dark pisces.20120229.10318.fits DARK
" " 8 of 20 10.0000 blank/dark pisces.20120229.10319.fits DARK
" " 9 of 20 10.0000 blank/dark pisces.20120229.10320.fits DARK
" " 10 of 20 10.0000 blank/dark pisces.20120229.10321.fits DARK
" " 11 of 20 10.0000 blank/dark pisces.20120229.10322.fits DARK
" " 12 of 20 10.0000 blank/dark pisces.20120229.10323.fits DARK
" " 13 of 20 10.0000 blank/dark pisces.20120229.10324.fits DARK
" " 14 of 20 10.0000 blank/dark pisces.20120229.10325.fits DARK
" " 15 of 20 10.0000 blank/dark pisces.20120229.10326.fits DARK
" " 16 of 20 10.0000 blank/dark pisces.20120229.10327.fits DARK
" " 17 of 20 10.0000 blank/dark pisces.20120229.10328.fits DARK
" " 18 of 20 10.0000 blank/dark pisces.20120229.10329.fits DARK
" " 19 of 20 10.0000 blank/dark pisces.20120229.10330.fits DARK
" " 20 of 20 10.0000 blank/dark pisces.20120229.10331.fits DARK
" " 1 of 20 25.0000 blank/dark pisces.20120229.10332.fits DARK
" " 2 of 20 25.0000 blank/dark pisces.20120229.10333.fits DARK
" " 3 of 20 25.0000 blank/dark pisces.20120229.10334.fits DARK
" " 4 of 20 25.0000 blank/dark pisces.20120229.10335.fits DARK
" " 5 of 20 25.0000 blank/dark pisces.20120229.10336.fits DARK
" " 6 of 20 25.0000 blank/dark pisces.20120229.10337.fits DARK
" " 7 of 20 25.0000 blank/dark pisces.20120229.10338.fits DARK
" " 8 of 20 25.0000 blank/dark pisces.20120229.10339.fits DARK
" " 9 of 20 25.0000 blank/dark pisces.20120229.10340.fits DARK
" " 10 of 20 25.0000 blank/dark pisces.20120229.10341.fits DARK
" " 11 of 20 25.0000 blank/dark pisces.20120229.10342.fits DARK
" " 12 of 20 25.0000 blank/dark pisces.20120229.10343.fits DARK
" " 13 of 20 25.0000 blank/dark pisces.20120229.10344.fits DARK
" " 14 of 20 25.0000 blank/dark pisces.20120229.10345.fits DARK
" " 15 of 20 25.0000 blank/dark pisces.20120229.10346.fits DARK
" " 16 of 20 25.0000 blank/dark pisces.20120229.10347.fits DARK
" " 17 of 20 25.0000 blank/dark pisces.20120229.10348.fits DARK
" " 18 of 20 25.0000 blank/dark pisces.20120229.10349.fits DARK
" " 19 of 20 25.0000 blank/dark pisces.20120229.10350.fits DARK
" " 20 of 20 25.0000 blank/dark pisces.20120229.10351.fits DARK

Close

14:27 Close dome T=-2 degC D=-26 degC Wind 10 m/sec from West. Sky remains photometric although the scale height of dust appears well above the mountaintop.

-- JohnHill - 29 Feb 2012
I Attachment Action Size Date Who Comment
AO423.obsobs AO423.obs manage 624 bytes 26 Mar 2012 - 22:43 JohnHill  
AO499.obsobs AO499.obs manage 690 bytes 26 Mar 2012 - 22:44 JohnHill  
BS9142_test.obsobs BS9142_test.obs manage 250 bytes 26 Mar 2012 - 22:37 JohnHill  
HIP43121_sat.obsobs HIP43121_sat.obs manage 309 bytes 26 Mar 2012 - 22:39 JohnHill  
HIP43121_test.obsobs HIP43121_test.obs manage 206 bytes 26 Mar 2012 - 22:38 JohnHill  
Q1135_presetH.obsobs Q1135_presetH.obs manage 956 bytes 26 Mar 2012 - 22:42 JohnHill  
c2009p1garradd_J.obsobs c2009p1garradd_J.obs manage 685 bytes 26 Mar 2012 - 22:46 JohnHill  
c2009p1garradd_J_new.obsobs c2009p1garradd_J_new.obs manage 437 bytes 26 Mar 2012 - 22:47 JohnHill  
c2009p1garradd_Ks.obsobs c2009p1garradd_Ks.obs manage 755 bytes 26 Mar 2012 - 22:44 JohnHill  
c2009p1garradd_bg.obsobs c2009p1garradd_bg.obs manage 314 bytes 26 Mar 2012 - 22:46 JohnHill  
c2009p1garradd_test.obsobs c2009p1garradd_test.obs manage 212 bytes 26 Mar 2012 - 22:40 JohnHill  
dgtau_test.obsobs dgtau_test.obs manage 218 bytes 26 Mar 2012 - 22:40 JohnHill  
Topic revision: r12 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback